Supposes-102 + 1.02-+ 1.025 +-+ answer to two decimal places. 1.02 1.02 + יהן . Obtain the value of S. Round your final 1.02 1.0
Answer #1

Checking for common ratio, r

T2/T1 = 1/(1.02)^3/1/(1.02)

= 1/(1.02)^2

Similarly, T3/T2 = 1/(1.02)^5/1/(1.02)^3

= 1/(1.02)^2

Since r is same, the given series is a geometric progression.

With a I.e. first term = 1/1.02

And r I.e. common ratio = 1/(1.02)^2

Sum of infinite GP = a/(1-r)

= (1/1.02)/(1-1/1.0404)

=0.980392156/0.038831218

= 25.25 (approx.)